Molander Corporation is a distributor of a sun umbrella used at resort hotels. Data concerning the next month’s budget appear below:

Selling price per unit $ 28

Variable expense per unit $ 17

Fixed expense per month $ 8,910

Unit sales per month 960

Required: 1. What is the company’s margin of safety? (Do not round intermediate calculations.)

2. What is the company’s margin of safety as a percentage of its sales? (Round your percentage answer to 2 decimal places (i.e. .1234 should be entered as 12.34).)
Answer #1

Step 1: Calculation of Break even Sales Break even Sales = Fixed Cost/Contribution margin ratio = 8910/(28-17)/28 =8910/39.29